Giving a test to a group of students, the grades and gender are summarized below
A B C Total
Male 4 17 20 41
Female 15 7 19 41
Total 19 24 39 82

If one student is chosen at random,
Find the probability that the student did NOT get an "B"

Answer by ewatrrr(24785)


Hi
Giving a test to a group of students, the grades and gender are summarized below
A B C Total
Male 4 17 20 41
Female 15 7 19 41
Total 19 24 39 82
P(NOT B) = 1 - P(B) = 1 - 24/82 = 58/82 = 29/41